OpenStudy (aonz):

I forgot how to factorise this! :( 3x^2 +2xy - 8y^2 - 8x +14y-3

OpenStudy (anonymous):

Is this your equation? \[3x^2 +2xy - 8y^2 - 8x +14y-3\]

OpenStudy (aonz):

yes

OpenStudy (anonymous):

Sorry, I was trying to get my question answered.. Anyways! What do we know? We know that the following can be factored out: \[x^2\]\[8y^2\]\[14\] I left out 3 and why because they can only be only be factored out by 1.

OpenStudy (anonymous):

Not "why". I meant "y"

OpenStudy (aonz):

umm ok...

OpenStudy (anonymous):

Am I not making any sense? damn

OpenStudy (aonz):

for some reason i cant factorise this question :/ i dont see taking out these common factors to be any help...

OpenStudy (aonz):

i know its possible to factorise this but i ahve no idea how http://www.wolframalpha.com/input/?i=factorize+3x%5E2+%2B2xy+-+8y%5E2+-+8x+%2B14y-3

OpenStudy (anonymous):

It will! Promise.

OpenStudy (aonz):

umm, what method of factorising is this?

OpenStudy (anonymous):

REVERSE FOIL Starting with 3x^2 and 8y^2, what can they be factored into?

OpenStudy (anonymous):

3 and 1, x and x 2 and 4, y and y

OpenStudy (aonz):

sorry, but i dont understand what your talking about :/

OpenStudy (anonymous):

\[3x \times x = 3x^2\]\[2y \times 4y = 8y^2\] Is that not correct?

OpenStudy (aonz):

yes thats correct

OpenStudy (anonymous):

But its a negative so \[-4y \times 2y = -8y^2\]

OpenStudy (aonz):

ok

OpenStudy (anonymous):

But really from that link you posted, you can take out \[(x+2y-3)\] From the this you can take out x \[3x^2 + 2xy - 8x\] From this you can factor out 2y \[2xy - 8y^2 +14y\] And lastly -3\[-3\]

OpenStudy (anonymous):

...."you can factor out"....

OpenStudy (anonymous):

I guess you can I tried to find the greatest common factor in each, instead of reverse foil. Reverse FOIL is mostly for trinomials..

OpenStudy (aonz):

mhm i tried that. Btw i never done reverse foil before so is there any other way?

OpenStudy (aonz):

is it possible if you find the greatest common factor?

OpenStudy (anonymous):

Yes. I just showed you... Also some methods are: - Number of Terms - Factor Out the GCF First - Reversing FOIL - Guess and Check

OpenStudy (anonymous):

Its possible because that is what I did. For your x, you can factor out x ONLY because 3 is not a factor of 8. Also the lowest variable is x in all three. \[3x^2+2xy−8x\]

OpenStudy (aonz):